What is the radius of a sphere with a volume of 113.1 in.
Answer by atique.ah(27)

Solution:
V0lume of the sphere=4/3pir^3
113.1=4/3*pi*r^3
r^3=(113.1*3)/(4*pi)
r={(113.1*3)/(4*pi)}^1/3
where pi=22/7
solve this by using a calculator
it is aproximatelt 3inches